hsueu3guewhdiwhfu4:

In the last three years, Frederico’s basketball team won 30 more games than they lost. If they won 140 games, what was their ratio of wins to losses? Show the ratio in three different ways.

Vocaloid:

the first step is to calculate how many games were won, and how many were lost. "won 30 more games than they lost" so, let's choose some variable (I'll choose L) to represent how many games were lost. likewise, I'll choose W for wins. translating "won 30 more games than they lost", W = L + 30 in total, wins + losses = 140 total games, so L + W = 140 if we plug in W = L + 30 for L into the above equation, we get L + (L + 30) = 140 solve for L. then, since W = L + 30, we can simply find W by adding 30 to L

Vocaloid:

now, for the actual ratios: there are three ways we can write them: 1. ___ to ___ in this case, since it's "wins to losses" put the # of wins in the first blank and losses in the second blank 2. with a colon (__:__) similarly, wins go in the first blank and losses go into the second blank 3. with a fraction (___/___) similarly, wins go in the numerator and losses go into the denominator